# Your Language Model Can Secretly Write Like Humans: Contrastive Paraphrase Attacks on LLM-Generated Text Detectors

### v2

@@ Abstract
The misuse of large language models (LLMs), such as academic plagiarism, has driven the development of detectors to identify LLM-generated texts. To bypass these detectors, paraphrase attacks have emerged to purposely rewrite these texts to evade detection. Despite the success, existing methods require substantial data and computational budgets to train a specialized paraphraser, and their attack efficacy greatly reduces when faced with advanced detection algorithms. To address this, we propose Contrastive Paraphrase Attack (CoPA), a training-free method that effectively deceives text detectors using off-the-shelf LLMs. The first step is to carefully craft instructions that encourage LLMs to produce more human-like texts. Nonetheless, we observe that the inherent statistical biases of LLMs can still result in some generated texts carrying certain machine-like attributes that can be captured by detectors. To overcome this, CoPA constructs an auxiliary machine-like word distribution as a contrast to the human-like distribution generated by the LLM. By subtracting the machine-like patterns from the human-like distribution during the decoding process, CoPA is able to produce sentences that are less discernible by text detectors. Our theoretical analysis suggests the superiority of the proposed attack. Extensive experiments validate the effectiveness of CoPA in fooling text detectors across various scenarios.
@@
Hao Fang1, Jiawei Kong1,2, Tianqu Zhuang1, Yixiang Qiu1, Kuofeng Gao1, Bin Chen2,3, Shu-Tao Xia1,3, Yaowei Wang2,3, Min Zhang2, 1Tsinghua Shenzhen International Graduate School, Tsinghua University, 2Harbin Institute of Technology, Shenzhen, 3Pengcheng Laboratory Equal contribution, Correspondence: chenbin2021@hit.edu.cn
@@ 1 Introduction
Large language models (LLMs), such as GPT-4 and Claude-3.5, have demonstrated remarkable abilities in text comprehension and coherent text generation. These capabilities have driven their widespread applications, including code generation Jiang et al. (2024) and academic research Stokel-Walker (2022). However, the misuse of LLMs for harmful purposes, such as academic plagiarism and misinformation generation, has raised significant societal concerns Bao et al. (2024) regarding safety and ethics. In response, various detection methods that leverage the unique characteristics of LLM-generated texts from multiple perspectives have been proposed to mitigate the associated risks.
@@ 2.1 LLM-generated Text Detection
Existing detection algorithms for AI-generated text can generally be categorized into two types: (i) Training-based detection: These methods typically involve training a binary classification language model. Specifically, OpenAI employs a RoBERTa model Liu (2019) trained on a collection of millions of texts for detection. To enhance the detection robustness, RADAR Hu et al. (2023) draws inspiration from GANs Goodfellow et al. (2020) and incorporates adversarial training between a paraphraser and a detector. Additionally, DeTective Guo et al. (2024) proposes a contrastive learning framework to train the encoder to distinguish various writing styles of texts, and combined with a pre-encoding embedding database for classification. R-detect Song et al. (2025) employs a kernel relative test to judge a text by determining whether its distribution is closer to that of human texts. Despite the efforts in specific domains, training-based methods are struggling with generalization to unseen language domains, which reduces their practicality and versatility. (ii) Zero-shot detection: These methods are training-free and typically focus on extracting inherent features of LLMs’ texts to make decisions. GLTR Gehrmann et al. (2019) and LogRankSolaiman et al. (2019) leverage the probability or rank of the next token for detection. Since AI-generated text typically exhibits a higher probability than its perturbed version, DetectGPT Mitchell et al. (2023) proposes the probability curvature to distinguish LLM and human-written text. Building on this, Fast-DetectGPT Bao et al. (2024) greatly improves the efficiency by introducing conditional probability curvature that substitutes the perturbation step with a more efficient sampling step. TOCSIN Ma and Wang (2024) presents a plug-and-play module that incorporates random token deletion and semantic difference measurement to bolster zero-shot detection capabilities. Other methods explore different characteristics, such as likelihood Hashimoto et al. (2019), N-gram divergence Yang et al. (2024), and the editing distance from the paraphrased version Mao et al. (2024).
@@ 3.1 Problem Formulation
3.2 Contrastive Paraphrase Atatcks
@@
A reasonable solution involves constructing $p_{m}$ as a highly salient and concentrated machine-style token distribution, wherein high-probability tokens are strongly machine-related and more likely to trigger detection. We achieve this by identifying a prompt $x_{m}$ that significantly amplifies the LLM likelihood of the generated sentences. Figure 3 empirically validates our strategy, i.e., magnifying machine-style features in $p_{m}$ can, in turn, promote a refined distribution that more closely resembles authentic human writing, further enhancing the attack on the machine-text detector.
Adaptive Truncation for plausibility. Another important issue is that CoPA utilizes the whole token distribution to measure the difference. However, there may be occasions where certain high-probability tokens overlap between $p_{h}^{\prime}$ and $p_{m}$. The subtraction operation may penalize the probabilities of reasonable and valid tokens while rewarding casual and unrelated ones, thus compromising the coherence and semantic consistency of generated sentences. To address this, we incorporate a token constraint mechanism Li et al. (2023), which applies an adaptive pruning to the output tokens:
@@ 3.3 Theoretical Analysis
Figure 4: Illustration of the premise of Theorem 1. Let $|\mathcal{V}|=3$ and thus $\mathbb{P}^{\mathcal{V}}$ is a triangle. We draw the contours of $f(p)\coloneqq\mathbb{KL}(p_{h}||p)$. The closer to $p_{h}$, the lower the KL divergence with $p_{h}$. If $p_{h}^{\prime}-p_{m}$ points to the inside of the contour at $p_{h}^{\prime}$, then $f(p)$ decreases at $p_{h}^{\prime}$ along $p_{h}^{\prime}-p_{m}$, i.e., $g(\lambda)$ decreases at $\lambda=0$. In this case $g^{\prime}(0)<0$ is satisfied and Theorem 1 is applicable. In practice $p_{h}^{\prime}$ is usually between $p_{m}$ and $p_{h}$, so $g^{\prime}(0)<0$ is usually satisfied and Theorem 1 is generally applicable.
Note that the premise of Theorem 1 is $g^{\prime}(0)<0$. We use $|\mathcal{V}|=3$ as an example to illustrate its rationality. As in Figure 4, we calculate the area wherein probability distributions satisfy $g(0)^{\prime}<0$. In essence, $p_{h}^{\prime}$ is generated by the LLM and hence is constrained by the inherent language priors, which prevents it from deviating significantly from the machine-featured distribution $p_{m}$. Meanwhile, we use a carefully crafted human-like prompt to guide $p_{h}^{\prime}$ to move from $p_{m}$ towards $p_{h}$. As a result, $p_{h}^{\prime}$ typically falls within the region that satisfies $g^{\prime}(0)<0$. Therefore, this premise generally holds and thus Theorem 1 is applicable in practice, which is further confirmed by experimental results in Sec. 4.2. Based on LLM’s prediction paradigm, we contrast the output logits in practice, achieving excellent performance in misleading detection models.
@@ 4.1 Experimental Setup
Baselines. We compare our method with the state-of-the-art (SOTA) surrogate-free paraphrase attack Dipper Krishna et al. (2023). We adopt the setup of 60 lexical diversity and 60 order diversity for Dipper to achieve its best performance. Additionally, we reproduce the attack introduced in Mao et al. (2024), which leverages an LLM with the query “Help me rephrase it in human style.” to rewrite machine texts (denoted as Raidar-A). For fairness, we use the same LLM for both our paraphraser and baselines. Note that we also reveal the superiority of CoPA over Shi et al. (2024) that relies on an extra surrogate model in Appendix C.
@@
Metrics. We analyze the performance using two key metrics. (1) Detection accuracy. In real-world applications, it is crucial to guarantee that human-written text should almost never be misclassified as machine-generated Krishna et al. (2023), i.e., satisfying a very low false positive rate (FPR). Hence, we follow Dipper and report the true positive rate (TPR) at a fixed FPR. Specifically, we set a relatively high FPR of 5% to significantly reveal the performance improvements. Please refer to Appendix C for results with FPR=1%. (2) Semantic similarity. The rewritten sentences should preserve the original semantics. Similar to Dipper, we employ the P-SP Wieting et al. (2022) model, a specialized embedding model trained on a filtered paraphrase corpus Wieting and Gimpel (2018), to measure the semantic discrepancy. We align with Dipper and consider the semantics being preserved if the P-SP score exceeds the average real-human paraphrase score of 0.76. Moreover, we provide more analysis including text perplexity, GPT-4 assisted and human evaluation in Appendix G.
Implementation Details. For attack hyperparameters, we set the contrast intensity $\lambda=0.5$ and the clipping factor $\alpha=1\text{e}^{-5}$. Unless stated otherwise, we employ a single paraphrasing iteration. We employ Qwen2.5-72B-Instruct Qwen (2024) as the paraphraser. Due to page limits, we provide results paraphrased by four more LLMs (e.g., DeepSeek R1-Distill-32B DeepSeek-AI (2025)) in Appendix C. More details about the human and machine-like prompts are in Appendix B.
@@ 5 Conclusion
Limitation
While our method avoids the overhead of training a dedicated paraphraser by leveraging an off-the-shelf LLM, the contrastive paraphrasing mechanism requires two forward passes to construct the contrastive token distribution, bringing additional latency during next-token prediction. This may limit the practicality of the proposed attack in real-time applications. Besides, although human evaluation results in Appendix G indicate that CoPA-paraphrased texts are preferred over those from Dipper, we did not systematically account for the detailed linguistic backgrounds of evaluators and may introduce bias. A more comprehensive human study is needed to validate the general quality of the output sentences. Finally, this work follows prior studies and focuses exclusively on English text. Extending the contrastive paraphrasing framework to other languages such as Chinese and Spanish would be valuable for its broader applicability.
@@
This paper presents a novel method aimed to advance the research field of LLM-generated text detection. Note that all experiments are conducted within controlled laboratory environments. We do not expect the proposed method to serve as a powerful tool for potential adversaries but to raise society’s broader awareness of the vulnerability of current AI-text detectors. Also, the exceptional attack performance highlights the practical limitations of current detectors. Researchers of the open-source community are encouraged to conduct stress tests on their detectors against the proposed attack, based on which future studies can develop more robust stronger detectors. Furthermore, we conduct a preliminary study to alleviate the proposed threat via an adaptive defense that adversarially trains a Roberta-based detector using texts paraphrased by CoPA in Appendix F.
@@ Appendix A Theorem and Proof
$[-1,0]\subseteq I$ so $I$ is non-empty. $I$ is the intersection of some closed intervals, so $I$ is also a closed interval. Note that $g(\lambda)$ is continuous, so $g(\lambda)$ has a minimum value.
Assume that $g(\lambda)$ has minimum points $\lambda_{1}$ and $\lambda_{2}$ with $\lambda_{1}<\lambda_{2}$. By the convexity, $g(\lambda)$ is constant on $[\lambda_{1},\lambda_{2}]$. Note that $g(\lambda)$ is an analytic function. By Corollary 1.2.6 in Krantz and Parks (2002), the constancy of $g(\lambda)$ on $[\lambda_{1},\lambda_{2}]$ implies the constancy on $I$. This contradicts that $g(\lambda)$ is not constant, so the minimum point of $g(\lambda)$ is unique. ∎
@@
Corollary 1.2.6 in Krantz and Parks (2002) applies to open intervals, and $g(\lambda)$ is continuous at the endpoints, so it also applies to our closed intervals, $[\lambda_{1},\lambda_{2}]$ and $I$.
@@ Appendix B Experimental Details
(1) For human-style distributions:
@@
Considering that the original LLM-generated sentence carries the richest machine-related characteristics that are most easily detected, we adopt a direct and effective strategy by making LLM repeat the input sentence to obtain the most machine-like token probabilities.

### v3

@@ Abstract
The misuse of large language models (LLMs), such as academic plagiarism, has driven the development of detectors to identify LLM-generated texts. To bypass these detectors, paraphrase attacks have emerged to purposely rewrite these texts to evade detection. Despite the success, existing methods require substantial data and computational budgets to train a specialized paraphraser, and their attack efficacy greatly reduces when faced with advanced detection algorithms. To address this, we propose Contrastive Paraphrase Attack (CoPA), a training-free method that effectively deceives text detectors using off-the-shelf LLMs. The first step is to carefully craft instructions that encourage LLMs to produce more human-like texts. Nonetheless, we observe that the inherent statistical biases of LLMs can still result in some generated texts carrying certain machine-like attributes that can be captured by detectors. To overcome this, CoPA constructs an auxiliary machine-like word distribution as a contrast to the human-like distribution generated by the LLM. By subtracting the machine-like patterns from the human-like distribution during the decoding process, CoPA is able to produce sentences that are less discernible by text detectors. Our theoretical analysis suggests the superiority of the proposed attack. Extensive experiments validate the effectiveness of CoPA in fooling text detectors across various scenarios. The code is available at: https://github.com/ffhibnese/CoPA_Contrastive_Paraphrase_Attacks
@@
Hao Fangthanks: Equal Contribution1, Jiawei Kong1,2, Tianqu Zhuang1, Yixiang Qiu1, Kuofeng Gao1, Bin Chenthanks: Corresponding Author2,3, Shu-Tao Xia1,3, Yaowei Wang2,3, Min Zhang2, 1Tsinghua Shenzhen International Graduate School, Tsinghua University, 2Harbin Institute of Technology, Shenzhen, 3Pengcheng Laboratory {fangh25, kjw25, zhuangtq23, qiu-yx24, gkf21}@mail.tsinghua.edu.cn, chenbin2021@hit.edu.cn, xiast@sz.tsinghua.edu.cn, wangyw@pcl.ac.cn, zhangmin2021@hitsz.edu.cn
@@ 1 Introduction
Large language models (LLMs), such as GPT-4 and Claude-3.5, have demonstrated remarkable abilities in text comprehension and coherent text generation. These capabilities have driven their widespread applications, including code generation Jiang et al. (2024) and academic research Stokel-Walker (2022). However, the misuse of LLMs for harmful purposes, such as academic plagiarism and misinformation generation, has raised significant societal concerns regarding safety and ethics Bao et al. (2024); Wu et al. (2025); Kong et al. (2025). In response, various detection methods that leverage the unique characteristics of LLM-generated texts from multiple perspectives have been proposed to mitigate the associated risks.
@@ 2.1 LLM-generated Text Detection
Existing detection algorithms for AI-generated text can generally be categorized into two types: (i) Training-based detection: These methods typically involve training a binary classification language model. Specifically, OpenAI employs a RoBERTa model Liu (2019) trained on a collection of millions of texts for detection. To enhance the detection robustness, RADAR Hu et al. (2023) draws inspiration from GANs Goodfellow et al. (2020) and incorporates adversarial training between a paraphraser and a detector. Additionally, DeTective Guo et al. (2024a) proposes a contrastive learning framework to train the encoder to distinguish various writing styles of texts, and combined with a pre-encoding embedding database for classification. R-detect Song et al. (2025) employs a kernel relative test to judge a text by determining whether its distribution is closer to that of human texts. Despite the efforts in specific domains, training-based methods are struggling with generalization to unseen language domains, which reduces their practicality and versatility. (ii) Zero-shot detection: These methods are training-free and typically focus on extracting inherent features of LLMs’ texts to make decisions. GLTR Gehrmann et al. (2019) and LogRankSolaiman et al. (2019) leverage the probability or rank of the next token for detection. Since AI-generated text typically exhibits a higher probability than its perturbed version, DetectGPT Mitchell et al. (2023) proposes the probability curvature to distinguish LLM and human-written text. Building on this, Fast-DetectGPT Bao et al. (2024) greatly improves the efficiency by introducing conditional probability curvature that substitutes the perturbation step with a more efficient sampling step. TOCSIN Ma and Wang (2024) presents a plug-and-play module that incorporates random token deletion and semantic difference measurement to bolster zero-shot detection capabilities. Other methods explore different characteristics, such as likelihood Hashimoto et al. (2019), N-gram divergence Yang et al. (2024), and the editing distance from the paraphrased version Mao et al. (2024).
@@ 3.1 Problem Formulation
3.2 Contrastive Paraphrase Attacks
@@
A reasonable solution involves constructing $p_{m}$ as a highly salient and concentrated machine-style token distribution, wherein high-probability tokens are strongly machine-related and more likely to trigger detection. We achieve this by identifying a prompt $x_{m}$ that can amplify the LLM likelihood of generated sentences. Fig. 3 empirically validates our strategy, i.e., magnifying machine-style features in $p_{m}$ can, in turn, promote a refined distribution that more closely resembles authentic human writing, further enhancing the attack effectiveness.
Adaptive Truncation for plausibility. Another important issue is that CoPA utilizes the whole token distribution to measure the difference. However, there may be occasions where certain high-probability tokens overlap between $p_{h}^{\prime}$ and $p_{m}$. The subtraction operation may penalize the probabilities of reasonable and valid tokens while rewarding casual and unrelated ones Fang et al. (2025), thus compromising the coherence and semantic consistency of generated sentences. To address this, we incorporate a token constraint mechanism Li et al. (2023), which applies an adaptive pruning to the output tokens:
@@ 3.3 Theoretical Analysis
Figure 4: Illustration of the premise of Theorem˜1. Let $|\mathcal{V}|=3$ and thus $\mathbb{P}^{\mathcal{V}}$ is a triangle. We draw the contours of $f(p)\coloneqq\mathbb{KL}(p_{h}||p)$. The closer to $p_{h}$, the lower the KL divergence with $p_{h}$. If $p_{h}^{\prime}-p_{m}$ points to the inside of the contour at $p_{h}^{\prime}$, then $f(p)$ decreases at $p_{h}^{\prime}$ along $p_{h}^{\prime}-p_{m}$, i.e., $g(\lambda)$ decreases at $\lambda=0$. In this case $g^{\prime}(0)<0$ is satisfied and Theorem˜1 is applicable. In practice $p_{h}^{\prime}$ is usually between $p_{m}$ and $p_{h}$, so $g^{\prime}(0)<0$ is usually satisfied and Theorem˜1 is generally applicable.
Note that the premise of Theorem˜1 is $g^{\prime}(0)<0$. We use $|\mathcal{V}|=3$ as an example to illustrate its rationality. As in Figure 4, we calculate the area wherein probability distributions satisfy $g(0)^{\prime}<0$. In essence, $p_{h}^{\prime}$ is generated by the LLM and hence is constrained by the inherent language priors, which prevent it from deviating significantly from the machine-featured distribution $p_{m}$. Meanwhile, we use a carefully crafted human-like prompt to guide $p_{h}^{\prime}$ to move from $p_{m}$ towards $p_{h}$. As a result, $p_{h}^{\prime}$ typically falls within the region that satisfies $g^{\prime}(0)<0$. Therefore, this premise generally holds and thus Theorem˜1 is applicable in practice, which is further confirmed by experimental results in Sec. 4.2. Based on LLM’s prediction paradigm, we contrast the output logits in practice, achieving excellent performance in misleading detection models. We also note that researchers should examine the validity of this assumption in their specific setting before applying our theoretical framework.
@@ 4.1 Experimental Setup
Baselines. We compare our method with the state-of-the-art (SOTA) surrogate-free paraphrase attack Dipper Krishna et al. (2023). We adopt the setup of 60 lexical diversity and 60 order diversity for Dipper to achieve its best performance. Additionally, we reproduce the attack introduced in Mao et al. (2024), which leverages an LLM with the query “Help me rephrase it in human style.” to rewrite machine texts (denoted as Raidar-A). For fairness, we use the same LLM for both our paraphraser and baselines. Note that we also reveal the superiority of CoPA over Shi et al. (2024) that relies on an extra surrogate model in Appendix˜C.
@@
Metrics. We analyze the performance using two key metrics. (1) Detection accuracy. In real-world applications, it is crucial to guarantee that human-written text should almost never be misclassified as machine-generated Krishna et al. (2023), i.e., satisfying a very low false positive rate (FPR). Hence, we follow Dipper and report the true positive rate (TPR) at a fixed FPR. Specifically, we set a relatively high FPR of 5% to significantly reveal the performance improvements. Please refer to Appendix C for results with FPR=1%. (2) Semantic similarity. The rewritten sentences should preserve the original semantics. Similar to Dipper, we employ the P-SP Wieting et al. (2022) model, a specialized embedding model trained on a filtered paraphrase corpus Wieting and Gimpel (2018), to measure the semantic discrepancy. We align with Dipper and consider the semantics being preserved if the P-SP score exceeds the average real-human paraphrase score of 0.76. Moreover, we provide more analysis including text perplexity, GPT-4 assisted and human evaluation in Appendix˜H.
Implementation Details. For attack hyperparameters, we set the contrast intensity $\lambda=0.5$ and the clipping factor $\alpha=1\text{e}^{-5}$. Unless stated otherwise, we employ a single paraphrasing iteration. We employ Qwen2.5-72B-Instruct Qwen (2024) as the paraphraser. Due to page limits, we provide results paraphrased by more LLMs in Appendix˜C. More details about the human and machine-like prompts are in Appendix B.
@@ 5 Conclusion
Limitations
While our method avoids the overhead of training a dedicated paraphraser by leveraging an off-the-shelf LLM, the contrastive paraphrasing mechanism requires two forward passes to construct the contrastive token distribution, bringing additional latency during next-token prediction. This may limit the practicality of the proposed attack in real-time applications. Moreover, using off-the-shelf LLMs to paraphrase texts can lead to the semantics of output texts deviating from the original sentence, which may require multiple generations to maintain the semantic similarity. Besides, although human evaluation results in Appendix H indicate that CoPA-paraphrased texts are preferred over those from Dipper, we did not systematically account for the detailed linguistic backgrounds of evaluators and may introduce bias. A more comprehensive human study is needed to validate the general quality of the output sentences. Finally, this work follows prior studies and focuses exclusively on English text. Extending the contrastive paraphrasing framework to other languages, such as Chinese and Spanish, would be valuable for its broader applicability.
@@
This paper presents a novel method aimed to advance the research field of LLM-generated text detection. Note that all experiments are conducted within controlled laboratory environments. We do not expect the proposed method to serve as a powerful tool for potential adversaries but to raise society’s broader awareness of the vulnerability of current AI-text detectors. Also, the exceptional attack performance highlights the practical limitations of current detectors. Researchers of the open-source community are encouraged to conduct stress tests on their detectors against the proposed attack, based on which future studies can develop more robust stronger detectors. Furthermore, we conduct a preliminary study to alleviate the proposed threat via an adaptive defense that adversarially trains a Roberta-based detector using texts paraphrased by CoPA in Appendix˜F.
@@
Acknowledgement
This work is supported in part by the National Natural Science Foundation of China under grant 62171248, 62301189, 62576122, and Shenzhen Science and Technology Program under Grant KJZD20240903103702004, JCYJ20220818101012025, GXWD20220811172936001.
@@ Appendix A Theorem and Proof
$1,0]\subseteq I$ so $I$ is non-empty. $I$ is the intersection of some closed intervals, so $I$ is also a closed interval. Note that $g(\lambda)$ is continuous, so $g(\lambda)$ has a minimum value.
Assume that $g(\lambda)$ has minimum points $\lambda_{1}$ and $\lambda_{2}$ with $\lambda_{1}<\lambda_{2}$. By the convexity, $g(\lambda)$ is constant on $\lambda_{1},\lambda_{2}]$. Note that $g(\lambda)$ is an analytic function. By Corollary 1.2.6 in Krantz and Parks (2002), the constancy of $g(\lambda)$ on $\lambda_{1},\lambda_{2}]$ implies the constancy on $I$. This contradicts that $g(\lambda)$ is not constant, so the minimum point of $g(\lambda)$ is unique. ∎
@@
Corollary 1.2.6 in Krantz and Parks (2002) applies to open intervals, and $g(\lambda)$ is continuous at the endpoints, so it also applies to our closed intervals, $\lambda_{1},\lambda_{2}]$ and $I$.
@@ Appendix B Experimental Details
(1) For human-style distributions: An intuitive way is to construct prompts that directly ask LLMs to produce human-like sentences by encouraging vivid lexical substitution and diverse sentence structures. However, compared to a direct request for human-toned texts, we empirically find that framing a realistic human conversation scene to LLMs leads to more vivid and diverse generations, strengthening the attack effectiveness.
@@
To elicit machine-like responses from LLMs, our prompt engineering (PE) constructs a wide range of prompts, e.g., ask for direct paraphrasing or instruct the LLM to reply in the tone of an AI assistant. The LLM likelihood provided by a detector is used as a proxy to quantify the degree of machine-style characteristics in the output. We find that naively prompting an LLM to rewrite a machine text often reduces existing machine-related features, as the rewritten texts inherit mixed stylistic and lexical preferences from multiple LLMs, increasing textual diversity. To preserve machine-related features, we incorporate stylistic cues into the prompts, e.g., instructions like "… in the tone of an AI assistant …" or "… in a machine-writting style …" to guide the LLM for a more typical machine expression, which indeed makes the output more easily detectable and, in turn, enhances the attack.
@@
Considering that the original LLM-generated sentence carries the richest machine-related characteristics that are most easily detected, we adopt a direct and effective strategy by making LLM repeat the input sentence to obtain the most machine-like token probabilities. We provide the results of representative prompts in Fig. 3, which validate the effectiveness of our selection strategy.